# PathFinder: Advancing Path Loss Prediction for Single-to-Multi-Transmitter Scenario

### v1

@@ Abstract
Radio path loss prediction (RPP) is critical for optimizing 5G networks and enabling IoT, smart city, and similar applications. However, current deep learning-based RPP methods lack proactive environmental modeling, struggle with realistic multi-transmitter scenarios, and generalize poorly under distribution shifts, particularly when training/testing environments differ in building density or transmitter configurations. This paper identifies three key issues: (1) passive environmental modeling that overlooks transmitters and key environmental features; (2) overemphasis on single-transmitter scenarios despite real-world multi-transmitter prevalence; (3) excessive focus on in-distribution performance while neglecting distribution shift challenges. To address these, we propose PathFinder, a novel architecture that actively models buildings and transmitters via disentangled feature encoding and integrates Mask-Guided Low-rank Attention to independently focus on receiver and building regions. We also introduce a Transmitter-Oriented Mixup strategy for robust training and a new benchmark, single-to-multi-transmitter RPP (S2MT-RPP), tailored to evaluate extrapolation performance (multi-transmitter testing after single-transmitter training). Experimental results show PathFinder outperforms state-of-the-art methods significantly, especially in challenging multi-transmitter scenarios. Our code and project site are available at: https://emorzz1g.github.io/PathFinder/
keywords:
Wireless Communication , Radio Path Loss Prediction , Distribution Shift , Multi-Transmitter , 5G Communication
journal: Pattern Recognition
[…]
[label1] organization=South China University of Technology, addressline=, city=Guangzhou, postcode=510650, state=Guangdong, country=China
[…]
[label2] organization=Pengcheng Laboratory, addressline=, city=Shenzhen, postcode=518066, state=Guangdong, country=China
1 Introduction
The rapid advancement of 5G communication highlights benefits like low latency and extensive coverage, establishing a fundamental infrastructure for applications in the Internet of Things, intelligent driving, and smart cities [bg_rpp, wifi]. Among these, radio path loss prediction is crucial in communication network planning, signal forecasting, and drone base station/transmitter deployment. In essence, the Radio Path loss Prediction (RPP) task aims to predict the information loss distribution across regions based on the base station location and relevant data within a specific building environment [rpp2, rpp3, RadioDIP]. This is shown in Fig. 1(a).
@@ 1 Introduction
The ray tracing method is widely used for predicting wireless point path loss in mobile communication environments [dataset, RadioUNet]. While this method provides accurate predictions, its intricate modeling and computational processes make it unsuitable for real-time applications [RadioUNet, ml_rpp1]. Therefore, it is better suited for design tasks like base station site selection that do not require immediate results. Radio path loss prediction essentially involves estimating or generating predictions.
@@
2 Related Work
2.1 Radio Pathloss Prediction
@@
2.2 Distribution Shift in Radio Path Loss Prediction
@@
3 Preliminaries
3.1 Problem Definition
@@ 4 PathFinder Model
4.1 Disentangled Feature Encoding
@@
4.2 Mask-Guided Low-rank Attention
@@
${\bm{R}}_{token}=(\operatorname{LN}\circ\operatorname{Reshape}\circ\mathcal{F}_{LR})({\bm{X}}_{1}),\ (\mathbb{R}^{H\times W\times D}\rightarrow\mathbb{R}^{HW\times E})$
${\bm{P}}_{token}=\operatorname{LN}({\bm{P}}),\ (\mathbb{R}^{n\times E}\rightarrow\mathbb{R}^{n\times E}),$
@@
${\bm{A}}_{B}=\operatorname{Fill}({\bm{A}}_{B},{\bm{M}}^{\prime})\in\mathbb{R}^{HW\times n},{\bm{A}}_{R}=\operatorname{Fill}({\bm{A}}_{R},{\bm{M}})\in\mathbb{R}^{HW\times n}.$ (7)
@@
${\bm{O}}_{2}=\operatorname{Softmax}({\bm{A}}_{R}){\bm{\mathsfit{V}}}_{P},\ (\mathbb{R}^{HW\times n}\rightarrow\mathbb{R}^{HW\times E})$
${\bm{O}}=\operatorname{Conv}\left(\frac{{\bm{O}}_{1}+{\bm{O}}_{2}}{2}{\bm{\mathsfit{W}}}_{O}\right),\ (\mathbb{R}^{HW\times E}\rightarrow\mathbb{R}^{HW\times E}),$
${\bm{O}}=\operatorname{Reshape}({\bm{O}})\ (\mathbb{R}^{HW\times E}\rightarrow\mathbb{R}^{H\times W\times D}),$
@@ 4.3 PathFinder Pipeline
4.4 Transmitter-Oriented Mixup
@@
4.5 Model Optimization
To ensure the accuracy of radio path loss map predictions, the mean squared error (MSE) loss is commonly employed to optimize the model. However, during the early stages of model training, when the model’s output significantly deviates from the target, the gradients of the MSE can become excessively large, leading to instability in training. Moreover, due to the substantial number of parameters in the UNet architecture, the slow optimization speed in the initial phase is exacerbated by the MSE loss. To address these shortcomings, we propose the Momentum Prediction Loss (MPL). This loss function introduces an additional parameter $\delta$ for adaptive momentum updates, smoothing the weights of the MSE loss and the mean absolute error (MAE) loss. It can be expressed as follows:
@@ 4.5 Model Optimization
5 Experiment
5.1 Experimental Setup
@@ 5.2 Comparison on DS-RPP
Table 1 shows various models’ performance on the testing set. Our model (PathFinder) outperforms others across both metrics, achieving SOTA results. It improves RMSE and RMSE-R metrics by 30.59% and 29.08% respectively, compared to REM-Net, significantly enhancing prediction performance in both global and local areas.
@@
Table 1: Comparison of different models’ performance DS-RPP. Bold indicates the best model, while underline indicates the second-best model.
@@ 5.3 Coverage Analysis
Table 2 presents all models’ performance regarding critical path loss area coverage. The final column shows our model’s relative improvement over the second best model. As the importance factor increases (from 40% to 5%), PathFinder’s advantages become more pronounced. For the 40% most important coverage area, performance improves by 34.46% over REM-Net. For only the 5% most important coverage area, PathFinderoutperforms REM-Net by 52.38%. This demonstrates our model significantly enhances path loss prediction accuracy across different importance ratios.
@@
Table 2: RMSE comparison of different models when addressing coverage areas of varying importance.
@@
5.4 Comparison on S2MT-RPP
To analyze model performance with multiple transmitters, we first established a DS-RPP task with two transmitters by fixing the parameter $\beta=0.5$ in Eq. (10) and omitting image enhancement. To further assess performance with varying transmitter counts, we modified Eq. (10) to average the effects of any number of ${\bm{S}}_{i}$ and ${\bm{Y}}_{i}$ for S2MT-RPP. We evaluated models with 2 to 5 transmitters, presenting RMSE results in Table 3 and visualizing cases with 2 and 4 transmitters in Figs. 6 and 7, respectively.
Table 3 shows that while baseline models like AE and RadioUNet experience increased loss and performance decline as transmitter count rises, more advanced models like PMNet and REM-Net maintain relatively consistent performance. Our model, however, demonstrates remarkable consistency across all transmitter counts, achieving a 59.35% improvement over REM-Net even with five transmitters. This highlights the effectiveness of the proposed TOM strategy in enhancing our model’s performance in distribution shift scenarios and its robust zero-shot capabilities.
@@ 5.4 Comparison on S2MT-RPP
Table 3: Comparison of different models’ performance under distribution shifts caused by varying numbers of transmitters.
@@
5.5 Convergence Analysis of Different Models
To further investigate the reasons behind the differences in distribution shift handling capabilities across various models and to demonstrate that the proposed MPL optimization effectively aids model convergence, we analyze the convergence processes of different models. Fig. 8 illustrates the changes in loss for all models during training and validation. "Ours w/o MPL" represents our model without MPL optimization, wherein MSE is used as a substitute loss for model optimization. In implementation, we set a maximum of 30 training epochs and employed an early stopping strategy, saving the best model when the validation loss no longer decreases. In our experiments, our model, along with PMNet and REM-Net, triggered this strategy.
@@ Acknowledgment
Appendix A Experiment
A.1 Evaluation Metric
@@ 3. Normalized Mean Squared Error for Receiver (NMSE-R):
A.2 Comparison on DS-RPP
Fig. 9 displays the prediction results of all baseline methods alongside our model across different samples.
A.3 Coverage Analysis
Figs. 10, 11, 12, 13 and 14 visualize the results of different models when the importance of the coverage area is 40%, 30%, 20%, 10% and 5%, respectively.
Figure 10: Visualization cases of different models when dealing with the most important 40% of the coverage area.
Figure 11: Visualization cases of different models when dealing with the most important 30% of the coverage area.
Figure 13: Visualization cases of different models when dealing with the most important 10% of the coverage area.
Figure 14: Visualization cases of different models when dealing with the most important 5% of the coverage area.
Table 4: Comparison of different models’ performance under distribution shifts caused by varying numbers of transmitters.

### v2

@@ Abstract
Radio path loss prediction (RPP) is critical for optimizing 5G networks and enabling IoT, smart city, and similar applications. However, current deep learning-based RPP methods lack proactive environmental modeling, struggle with realistic multi-transmitter scenarios, and generalize poorly under distribution shifts, particularly when training/testing environments differ in building density or transmitter configurations. This paper identifies three key issues: (1) passive environmental modeling that overlooks transmitters and key environmental features; (2) overemphasis on single-transmitter scenarios despite real-world multi-transmitter prevalence; (3) excessive focus on in-distribution performance while neglecting distribution shift challenges. To address these, we propose PathFinder, a novel architecture that actively models buildings and transmitters via disentangled feature encoding and integrates Mask-Guided Low-rank Attention to independently focus on receiver and building regions. We also introduce a Transmitter-Oriented Mixup strategy for robust training and a new benchmark, single-to-multi-transmitter RPP (S2MT-RPP), tailored to evaluate extrapolation performance (multi-transmitter testing after single-transmitter training). Experimental results show PathFinder outperforms state-of-the-art methods significantly, especially in challenging multi-transmitter scenarios. Our code and project site are publicly available [^1] https://emorzz1g.github.io/PathFinder/.
I Introduction
The rapid advancement of 5G communication highlights benefits like low latency and extensive coverage, establishing a fundamental infrastructure for applications in the Internet of Things, intelligent driving, and smart cities [bg_rpp, wifi, tmc_radiomap1, tmc_radiomap2]. Among these, radio path loss prediction is crucial in communication network planning, signal forecasting, and drone base station/transmitter deployment. In essence, the Radio Path loss Prediction (RPP) task aims to predict the information loss distribution across regions based on the base station location and relevant data within a specific building environment [rpp2, rpp3, tmc_radiomap3, tmc_radiomap4]. This is shown in Fig. 1(a).
@@ I Introduction
The ray tracing method is widely used for predicting radio path loss in mobile communication environments [dataset, RadioUNet]. While this method provides accurate predictions, its intricate modeling and computational processes make it unsuitable for real-time applications [RadioUNet, ml_rpp1]. Therefore, it is better suited for design tasks like base station site selection that do not require immediate results. Radio path loss prediction essentially involves estimating or generating predictions.
@@
II Related Work
II-A Radio Pathloss Prediction
@@
II-B Distribution Shift in Radio Path Loss Prediction
@@
III Preliminaries
III-A Problem Definition
@@ IV PathFinder Model
IV-A Disentangled Feature Encoding
@@
IV-B Mask-Guided Low-rank Attention
@@
${\bm{R}}_{token}=(\operatorname{LN}\circ\operatorname{Reshape}\circ\mathcal{F}_{LR})({\bm{X}}_{1}),\ (\mathbb{R}^{H\times W\times D}\rightarrow\mathbb{R}^{HW\times E})$ (6)
${\bm{P}}_{token}=\operatorname{LN}({\bm{P}}),\ (\mathbb{R}^{n\times E}\rightarrow\mathbb{R}^{n\times E}),$ (7)
@@
${\bm{A}}_{B}=\operatorname{Fill}({\bm{A}}_{B},{\bm{M}}^{\prime})\in\mathbb{R}^{HW\times n},$ (9)
${\bm{A}}_{R}=\operatorname{Fill}({\bm{A}}_{R},{\bm{M}})\in\mathbb{R}^{HW\times n}.$ (10)
@@
${\bm{O}}_{2}=\operatorname{Softmax}({\bm{A}}_{R}){\bm{\mathsfit{V}}}_{P},\ (\mathbb{R}^{HW\times n}\rightarrow\mathbb{R}^{HW\times E})$ (12)
${\bm{O}}=\operatorname{Conv}\left(\frac{{\bm{O}}_{1}+{\bm{O}}_{2}}{2}{\bm{\mathsfit{W}}}_{O}\right),\ (\mathbb{R}^{HW\times E}\rightarrow\mathbb{R}^{HW\times E}),$ (13)
${\bm{O}}=\operatorname{Reshape}({\bm{O}})\ (\mathbb{R}^{HW\times E}\rightarrow\mathbb{R}^{H\times W\times D}),$ (14)
@@ IV-C PathFinder Pipeline
IV-D Transmitter-Oriented Mixup
@@
IV-E Model Optimization
To ensure the accuracy of radio path loss map predictions, the mean squared error (MSE) loss [cpatchbls, adamembls] is commonly employed to optimize the model. However, during the early stages of model training, when the model’s output significantly deviates from the target, the gradients of the MSE can become excessively large, leading to instability in training. Moreover, due to the substantial number of parameters in the UNet architecture, the slow optimization speed in the initial phase is exacerbated by the MSE loss. To address these shortcomings, we propose the Momentum Prediction Loss (MPL). This loss function introduces an additional parameter $\delta$ for adaptive momentum updates, smoothing the weights of the MSE loss and the mean absolute error (MAE) loss. It can be expressed as follows:
@@ IV-E Model Optimization
V Theoretical Analysis
V-A Complexity Analysis of Mask-Guided Low-rank Attention
Standard self-attention mechanisms typically incur a quadratic computational complexity with respect to the input sequence length [simad, simad_conf]. Given a feature map with spatial dimensions $H\times W$, the sequence length is $N=HW$. The complexity of standard attention is $\mathcal{O}(N^{2}d)$, where $d$ is the feature dimension. This is computationally prohibitive for high-resolution path loss maps.
In contrast, our proposed Mask-Guided Low-rank Attention (MLA) significantly reduces this cost. By disentangling the transmitter prompt $\bm{P}$} from the environmental features, we utilize $\bm{P}$} as the bridge for attention computation. Let $n$ denote the number of transmitters (where $n\ll N$) and $E$ be the embedding dimension. The computation of queries ${\bm{Q}}_{B},{\bm{Q}}_{R}\in\mathbb{R}^{N\times E}$ and keys/values ${\bm{K}}_{P},{\bm{V}}_{P}\in\mathbb{R}^{n\times E}$ involves linear projections with complexity $\mathcal{O}(NE^{2}+nE^{2})$.
The cross-attention score calculation, e.g., ${\bm{A}}_{B}={\bm{Q}}_{B}{\bm{K}}_{P}^{\top}$, involves the multiplication of an $(N\times E)$ matrix by an $(E\times n)$ matrix, resulting in a complexity of $\mathcal{O}(NnE)$. Similarly, the aggregation step ${\bm{O}}_{1}=\text{Softmax}({\bm{A}}_{B}){\bm{V}}_{P}$ incurs $\mathcal{O}(NnE)$. Since $n$ is a small constant (representing the number of transmitter prompts) and $E$ is the low-rank dimension, the total complexity of our MLA is:
$\mathcal{C}_{MLA}=\mathcal{O}(NnE)\approx\mathcal{O}(N).$ (18)
Thus, PathFinder  achieves $\mathcal{O}(N)$ complexity with respect to the image resolution, ensuring scalability for large-scale urban environments compared to the $\mathcal{O}(N^{2})$ complexity of traditional methods.
V-B Theoretical Justification for S2MT Generalization
The core challenge of the S2MT-RPP task is extrapolating from a domain $\mathcal{D}_{single}$ to $\mathcal{D}_{multi}$. We justify the effectiveness of our Transmitter-Oriented Mixup (TOM) based on the physical principle of signal superposition.
V-B1 Physical Superposition Consistency
In electromagnetic theory, the received signal strength from multiple incoherent sources at a location $\mathbf{r}$ approximates the sum of the powers from individual sources (in linear scale). Let $\Phi({\bm{X}},{\bm{S}})$ be the ideal physical propagation function. For a set of transmitters ${\mathbb{S}}=\{{\textnormal{s}}_{1},\dots,{\textnormal{s}}_{k}\}$, the total path loss map satisfies the additivity property locally:
$\Phi({\bm{X}},\sum_{i=1}^{k}{\bm{S}}_{i})\approx\sum_{i=1}^{k}\Phi({\bm{X}},{\bm{S}}_{i}).$ (19)
Standard training on single-transmitter data $\mathcal{D}_{single}$ does not guarantee that the neural network $f_{\theta}$ satisfies Eq. (19).
V-B2 Linearity Inductive Bias via TOM
Our Transmitter-Oriented Mixup strategy constructs synthetic training samples via convex combinations. By minimizing the loss on mixed samples:
$\mathcal{L}_{mix}=\|f_{\theta}({\bm{X}},\beta{\bm{S}}_{i}+(1-\beta){\bm{S}}_{j})-(\beta{\bm{Y}}_{i}+(1-\beta){\bm{Y}}_{j})\|^{2},$ (20)
we explicitly regularize the model $f_{\theta}$ to behave linearly with respect to the transmitter input space $\bm{S}$}. Let the linearization error be $\epsilon_{lin}=|f_{\theta}(\lambda{\bm{S}}_{1}+\mu{\bm{S}}_{2})-(\lambda f_{\theta}({\bm{S}}_{1})+\mu f_{\theta}({\bm{S}}_{2}))|$. The TOM strategy minimizes the upper bound of $\epsilon_{lin}$. Consequently, when testing on a multi-transmitter input ${\bm{S}}_{test}=\sum{\bm{S}}^{(i)}$, the model prediction approximates the sum of individual predictions:
$f_{\theta}({\bm{X}},{\bm{S}}_{test})\approx\sum f_{\theta}({\bm{X}},{\bm{S}}^{(i)}).$ (21)
This theoretical property ensures that PathFinder  can extrapolate to $N$-transmitter scenarios even when trained only on single-transmitter data, effectively addressing the distribution shift problem defined in Problem 2.
V-C Gradient Smoothness of Momentum Prediction Loss
The proposed Momentum Prediction Loss (MPL) dynamically transitions between $L_{2}$ (MSE) and $L_{1}$ (MAE) loss based on an adaptive threshold $\delta$. Let $r={\bm{Y}}-\hat{{\bm{Y}}}$ be the residual. The gradient of MPL w.r.t prediction $\hat{{\bm{Y}}}$ is:
$\frac{\partial\mathcal{L}}{\partial\hat{{\bm{Y}}}}=\begin{cases}-r,&|r|\leq\delta\\ -\delta\cdot\text{sign}(r),&|r|>\delta\end{cases}$ (22)
Unlike standard Huber loss with a fixed $\delta$, our momentum-updated $\delta_{t}$ adapts to the training dynamics. In the early stages, residuals $|r|$ are large. Since $\delta_{t}$ tracks the Mean Absolute Error (MAE), the gradient is clipped to $\delta_{t}$, preventing exploding gradients commonly seen in MSE (where $\nabla\propto r$). As training converges and $|r|\to 0$, $\delta_{t}$ decreases, and the loss behaves like MSE, providing smooth gradients for fine-grained convergence. This adaptive mechanism ensures robust optimization across varying scales of path loss values.
VI Experiment
VI-A Experimental Setup
@@ VI-B Comparison on DS-RPP
Table I shows various models’ performance on the testing set. Our model (PathFinder) outperforms others across both metrics, achieving SOTA results. It improves RMSE and RMSE-R metrics by 30.59% and 29.08% respectively, compared to REM-Net, significantly enhancing prediction performance in both global and local areas.
@@
TABLE I: Comparison of different models’ performance DS-RPP. Bold indicates the best model, while underline indicates the second-best model.
@@ VI-C Coverage Analysis
Table II presents all models’ performance regarding critical path loss area coverage. The final column shows our model’s relative improvement over the second best model. As the importance factor increases (from 40% to 5%), PathFinder’s advantages become more pronounced. For the 40% most important coverage area, performance improves by 34.46% over REM-Net. For only the 5% most important coverage area, PathFinder  outperforms REM-Net by 52.38%. This demonstrates our model significantly enhances path loss prediction accuracy across different importance ratios.
@@
TABLE II: RMSE comparison of different models when addressing coverage areas of varying importance.
@@
VI-D Comparison on S2MT-RPP
To analyze model performance with multiple transmitters, we first established a DS-RPP task with two transmitters by fixing the parameter $\beta=0.5$ in Eq. (16) and omitting image enhancement. To further assess performance with varying transmitter counts, we modified Eq. (16) to average the effects of any number of ${\bm{S}}_{i}$ and ${\bm{Y}}_{i}$ for S2MT-RPP. We evaluated models with 2 to 5 transmitters, presenting RMSE results in Table III and visualizing cases with 2 and 4 transmitters in Figs. 6 and 7, respectively.
Table III shows that while baseline models like AE and RadioUNet experience increased loss and performance decline as transmitter count rises, more advanced models like PMNet and REM-Net maintain relatively consistent performance. Our model, however, demonstrates remarkable consistency across all transmitter counts, achieving a 59.35% improvement over REM-Net even with five transmitters. This highlights the effectiveness of the proposed TOM strategy in enhancing our model’s performance in distribution shift scenarios and its robust zero-shot capabilities.
@@ VI-D Comparison on S2MT-RPP
TABLE III: Comparison of different models’ performance under distribution shifts caused by varying numbers of transmitters.
@@
VI-E Convergence Analysis of Different Models
To further investigate the reasons behind the differences in distribution shift handling capabilities across various models and to demonstrate that the proposed MPL optimization effectively aids model convergence, we analyze the convergence processes of different models. Fig. 8 illustrates the changes in loss for all models during training and validation. ”Ours w/o MPL” represents our model without MPL optimization, wherein MSE is used as a substitute loss for model optimization. In implementation, we set a maximum of 30 training epochs and employed an early stopping strategy, saving the best model when the validation loss no longer decreases. In our experiments, our model, along with PMNet and REM-Net, triggered this strategy.
@@ Acknowledgments
A Experiment
A-A Evaluation Metric
@@ 3. Normalized Mean Squared Error for Receiver (NMSE-R):
A-B Comparison on DS-RPP
Fig. A9 displays the prediction results of all baseline methods alongside our model across different samples.
Figure A9: Visualization examples of different models in the DS-RPP task.
A-C Coverage Analysis
Figs. A10, A11, A12, A13 and A14 visualize the results of different models when the importance of the coverage area is 40%, 30%, 20%, 10% and 5%, respectively.
Figure A10: Visualization cases of different models when dealing with the most important 40% of the coverage area.
Figure A11: Visualization cases of different models when dealing with the most important 30% of the coverage area.
Figure A12: Visualization cases of different models when dealing with the most important 20% of the coverage area.
Figure A13: Visualization cases of different models when dealing with the most important 10% of the coverage area.
Figure A14: Visualization cases of different models when dealing with the most important 5% of the coverage area.
TABLE A4: Comparison of different models’ performance under distribution shifts caused by varying numbers of transmitters.